Consider a small economy that is closed to trade, so its net exports are equal to zero. Suppose that the economy has the following consumption function, where C is consumption, Y is real GDP, I is investment, G is government purchases, and T

is for net taxes:

C
=
30+0.75×(YT)

Suppose G

= $25 billion, I = $60 billion, and T

= $20 billion.

Given the consumption function and the fact that, in a closed economy, total expenditure can be calculated as Y=C+I+G

, the equilibrium output level isbillion.

Suppose the government purchases are increased by $50 billion. The new equilibrium level of output will be equal tobillion.

Based on the effect of the change in government purchases on equilibrium output, you can tell that this economy's spending multiplier is equal to   .
Answer #1

Answer:

1]

In equilibrium, Y = C + I + G

Y = 30 + 0.75(Y - 20) + 60 + 25

Y = 30 + 0.75Y – 15 + 85

Y = 100 + 0.75Y

0.25Y = 100

Y = $400 billion

2]

When G increases by 50, new level of G = 25 + 50 = 75

Y = 30 + 0.75(Y - 20) + 60 + 75

Y = 30 + 0.75Y – 15 + 135

0.25Y = 150

Y = $600 billion

3]

Spending multiplier = Increase in Y / Increase in G = $(600 - 400) Billion / $50 billion = 200/50 = 4